Feder Observatory

A Blog for Minnesota State University Moorhead's Paul P. Feder Observatory

Archive for the ‘Uncategorized’


Observing Log Oct. 13 0

Posted on October 14, 2016 by Erin Aadland

Tyler, Stefan, and Erin went out to the observatory with the intention of observing WASP-75 b.  The temperature was around 40 F, the focus was set to 2842, and the FWHM was between 4 and 7.  When we turned on the continuous shooting and the subframe  in order to focus the telescope, MaxIm DL decided it didn’t want to stop taking images. We had to focus the camera multiple times before it worked.  Then we tried observing WASP-75 b before realizing that the object was too close to the moon.  We also tried to get images of M31 but couldn’t get the exposure time right.

 

Observing Log 2/26/16 0

Posted on February 27, 2016 by Erin Aadland

Elias Holte, Laura Maixner, and I came out to observe M3.  The night started out with a camera error, where the camera wouldn’t connect to the computer.   After unplugging and replugging various cords, the camera finally connected.  Then after taking images of M3, the filter wheel had the error “Error=4” which was fixed by disconnecting the filter wheel in the MaxiumDL Setup tap of the camera control window.  We finished the night by taking pretty pictures of M51.

Temp = 38F

Focus = 2825

FWHM = 4.5

Observing Log 11/01/15 0

Posted on November 01, 2015 by Erin Aadland

Observed some pretty pictures of M15 tonight with Erin Aadland and Andy Block. When going through the procedure while setting up the filter wheel we kept getting “Error 1.”  Turned out the internet cable to the filter wheel was unplugged.  Once we fixed that, things went smoothly until the end.  Maxim DL shut down and when we tried to reconnect the telescope, we got an error reading “unable to connect.”  We couldn’t get it to work and since we only had two images left, we manually moved the telescope and called it a night.

Focus: 2840

Temperature: 50F

Log written by: Meghan Young but had to use Erin’s account 🙂

September 27, 2015 Observing Log 0

Posted on September 27, 2015 by Erin Aadland

So for once everything went according to plan.  Stefan Nelson, Ricquel Ramsbottle, and  Erin Aadland went observing during the lunar eclipse.  The night was full of eclipse watching and TrES-3b observing.

Focus:2879

Temp: 76F

September 12, 2015 Observing Log 0

Posted on September 13, 2015 by Erin Aadland

Andy Block, Erin Aadland, Stefan Nelson, and Jane Glanzer arrived at the telescope and got off to an optimistic start by taking the calibration images first.  However, disaster struck when the science images were about to commence.  First, the dome would not align with the telescope.  In fact, the dome faced the exact opposite direction as the dome.  The problem was resolved by manually moving the dome.  Second, as the first image was about to be taken Maxlm DL quit unexpectedly.  Third, an error occurred as we tried to reconnect the telescope.  We replugged the telescope and the problem was solved.  Fourth, the subframe images used to get a clear focus would not center on the star as needed.  A call from Matt Craig helped get around the issue by looking at the focus in the TCS window and looking at the FWHM of the star using the information window.  Finally, the science could commence (around 12:30am).  We ended up taking a few images of globular clusters for Laura Maixner and called it a night.

Temp: 69.1

Focus: 2889

August 30, 2015 Observing Log 0

Posted on August 30, 2015 by Erin Aadland

Erin Aadland, Andy Block, Jane Glanzer, and Dr. Craig went out to the Observatory to observe the effects of the smoke from South Texas on the TrES-3b star field.

Focus: 2889

Outside Temperature: 78 F

October 23, 2014 Observing Log 0

Posted on October 24, 2014 by Juan

Lauren U., Erin A., Brandon C., Brian T., and Eli H. went out to the telescope in hopes of observing an exo-planet transit, however, time ran short. So the M-29 cluster was observed instead. Skies were clear and the CCD camera actually worked. However, the computer froze when trying to get flats, darks, and biases, and had to be rebooted.

Observing log, Mar 1, 2013–Filter wheel is out of commission 0

Posted on March 01, 2013 by Matt

I came out tonight–the first clear night in almost two months–to observe some variable stars and take several images of standard star fields.

Unfortunately, the filter wheel wouldn’t work (ERROR=1) even after replacing the O-ring and cleaning the new O-ring.

The technical manual suggests there can be a problem with grounding if some of the RJ45 (ethernet) cable pins into the controller are bent, but they do not seem to be.

Next step will be to call Optec support and see what they say, and/or open up the filter housing and clean the motor shaft.

But, for now at least, the telescope is not usable.



↑ Top